OpenStudy (anonymous):

A sporting goods store has added a section with trail foods and snacks. Raisins cost $5.00 per pound and peanuts cost $3.80 per pound. Raisins and peanuts are going to be sold in a mixture of 30 pounds which will cost $4.05 per pound? How many pounds of raisins will be in the mix?

OpenStudy (anonymous):

Ultimately you will have a certain amount of raisins and peanuts that will add up to 30 lbs. Using this you can write your first equation. r+p=30 Now you know that mixing a certain amount of raisins and peanuts at their specified price will cost you $4.05. This leads to your second equation. 5.00(r) + 3.80(p) = 4.05 You have two equations with two variables now. Solve the first for one variable r = 30 - p plug into second equation 5.00(30 - p) + 3.80(p) = 4.05 Solve for p and you should be able to do the rest.

OpenStudy (anonymous):

do you get it?